# FakeFormer: Efficient Vulnerability-Driven Transformers for Generalisable Deepfake Detection

### v1

@@ 3 Vision Transformer and Deepfake Detection: Where is the Gap?
Performance with respect to the quality of deepfakes. Here, our goal is to quantify the performance of ViTs and CNNs with respect to the quality of the encountered deepfakes. To that aim, we compare in Figure 3(a) the performance of CNN-based methods (LAA-Net [45], CADDM [21], and SBI [47]) to transformer-based approaches (Plain ViT [23] and Swin [40]) [^2] Note that the small(-S) architecture configuration is employed for both ViT and Swin. The FLOPs and the number of parameters are reported in Section 5. across different Mask Self-Similarity Index Measure (SSIM) [42] ranges on the Celeb-DF [38] dataset. The Mask SSIM has been commonly used for assessing the quality of deepfakes [38, 55, 45] and is obtained by calculating the SSIM [59] score in the facial region between the fake and the corresponding original images. Higher Mask SSIM values indicate a higher perceptual quality of deepfakes [38, 55, 5, 45], generally implying the presence of more localized artifacts. It can be noted from Figure 3(a) that the performance of ViT drops more significantly for higher Mask SSIM values as compared to most CNN-based methods (LAA-Net and CADDM). The performance of Swin, on the other hand, does not deteriorate as much due to its ability to extract low-level features with its local window design; however, it still shows less stability than LAA-Net. These observations support our hypothesis. We specify that for a fair comparison with the state-of-the-art (SOTA), we train ViT and Swin using the same data synthesis [47] used in SBI and LAA-Net.
@@ 4.3.3 Learning-based Local Attention Module (L2-Att)
By predicting the locations of vulnerable patches, L2-Att diffuses artifact-prone local information through the ViT. In particular, L2-Att first takes the patch embeddings $\mathbf{z}^{L^{\prime}}_{\neg 0}\in\mathbb{R}^{N\times D}$ as input and processes them to produce spatial features as follows,
@@
Method Fake Test set (%)
CDF1 CDF2 DFW DFD DFDCP DFDC
@@
Xception [46] $\checkmark$ 58.81 65.59 61.18 66.93 65.29 55.37 89.75 85.48 69.90 91.98 45.60 -
FaceXRay [37] $\checkmark$ 80.58 73.33 - - - - 95.40 93.34 80.92 72.65 - -
LocalRL [13] $\checkmark$ - - 78.26 - - - 89.24 - 76.53 - - -
Multi-attentional [64] $\checkmark$ 69.14 74.03 68.26 75.25 73.56 73.79 92.95 96.51 83.81 96.52 - -
RECCE [7] $\checkmark$ 49.96 63.04 70.93 70.35 68.16 54.41 98.26 79.42 80.93 92.76 - -
SFDG [57] $\checkmark$ - - 75.83 - 69.27 - 88.00 - 73.63 - - -
CADDM [21] $\checkmark$ 89.36 93.25 93.88 91.12 74.48 75.23 99.03 99.59 - - 73.85 -
UCF [62] $\checkmark$ - - 82.4 - - - 94.5 - 80.5 - - -
M2TR [55] $\checkmark$ - - 68.2 - - - - - - - - -
DFDT [33] $\checkmark$ - - 88.3 - - - - - 76.1 - - -
TALL-Swin [60] $\checkmark$ - - 90.79 - - - - - 76.78 - - -
ICT [22] 81.43 - 85.71 - - - 84.13 - - - - -
SBI [47] 92.53 79.91 93.18 85.16 67.47 55.87 97.56 92.79 86.15 93.24 72.42 -
FakeFormer 97.25 98.36 94.45 97.15 81.74 83.72 96.12 98.31 96.30 99.50 78.91 80.01
Table 3: Cross-dataset evaluation. Comparisons in terms of AUC (%) and AP (%) on CDF1 [38], CDF2 [38], DFW [65], DFD [24], DFDCP [20], and DFDC [19]. ✓symbol is used to depict methods that utilized both Real data and Fake data for training.
@@ 5.1 Setups
Method Params Test set AUC(%)
@@
Method Fake Saturation Block Noise Blur Avg
FaceXray [37] $\checkmark$ 97.6 99.1 49.8 63.8 77.6
LipForensics [28] $\checkmark$ 99.9 87.4 73.8 96.1 89.3
LAA-Net [45] 99.9 99.9 53.9 98.2 87.9
FakeFormer 98.1 97.0 75.3 87.4 89.5
FakeSwin 99.8 99.9 81.6 89.8 92.8
Table 6: Robustness to perturbations on FF++ . Focusing on vulnerable patches, FakeFormer shows improved robustness to common perturbations, especially to noise.
@@ 5.2 Comparison with State-of-the-art
Blending-based Synthesis Approaches. Table 4 and Table 5 compare our method to SOTA when using Blending-Images (BI) [37] and Self-Blended-Images (SBI) [47], respectively. As observed, we generally achieve better performances than SOTA methods using the same blending algorithm on CDF2, DFDC, and DFDCP datasets, despite training on a drastically lower number of data as compared to other ViT-based approaches such as ICT [22]. Moreover, LAA-Net [45], as a CNN-based method, requires a larger model size as compared to our lightweight FakeFormer.
@@
Robustness to Perturbations. Since deepfakes can be altered by common perturbations on social media. Following the settings in [45, 28], experiments evaluating the robustness of FakeFormer against perturbations on FF++ [46] are reported in Table 6. We can observe the superiority of FakeFormer and FakeSwin as compared to existing methods.
@@ 5.3 Ablation Study
Parameter $\lambda$. The weight $\lambda$ defined in Eq. (9) is set empirically to $10$ as it yields the best performance on average. We report the results using different values of $\lambda$ within FakeFormer in Table 9. It can be observed the generalization across four testing benchmarks remains robust regardless of the value of $\lambda$.
@@ 8.2 More Details regarding the Datasets
Datasets. FF++ contains $1000$ original videos and $4000$ deepfake videos. The types of deepfakes include Deepfakes (DF) [48], Face2Face (F2F) [52], FaceSwap (FS) [35], and NeuralTextures (NT) [51]. Moreover, while CDF and DFD represent datasets including high-quality deepfakes, DFW stands out as one of the most challenging benchmarks. DFW is collected entirely from the internet without any prior knowledge regarding the used manipulation techniques. Additionally, it includes images with occlusions, illumination changes, different head poses, and diverse backgrounds. Meanwhile, DFDC and its preview version called DFDCP, comprise more than $128,000$ videos and is therefore one of the largest datasets in the field of deepfake detection.
Data Processing. The RetinaFace [18] algorithm is employed to extract facial bounding boxes. These boxes are enlarged by a factor of $1.25$ and then resized to a specific resolution $(H,W)$ to match the different backbone configurations. To generate pseudo-fakes, the two images are first aligned. For this purpose, the Dlib [34] library is utilized to extract $68$ and $81$ facial landmark for the two considered blending synthesis algorithms, BI [37] and SBI [47], respectively. Finally, the processed pseudo-fakes are generated for each training iteration.
Architecture Details. Here, we describe the hyperparameters of two considered FakeFormer variants:
@@
where the MLP size represents the dimension of hidden layers in MLP, the No. Heads denotes the number of heads in MHSA, Params is the number of parameters, and FLOPs represents the computational cost in terms of floating point operations per second. And for FakeSwin architecture, we adopt these two backbone variants from Swin [40], namely:
@@
where $M$ is the window size, $d$ is the query dimension of each head, the expansion layer of each MLP is $\alpha$, and $C_{h}$ denotes the number channel of the hidden layers in the first stage.
8.3 More Details of training Setups in Section 3
In this section, we present more details of the experiment settings in Section 3.
Specifically, in Figure 3a, we train vanilla ViT [23] and Swin [40] with SBI data synthesis [47] using the same training settings described in [45, 47]. Besides, the detection performance of three CNN methods in different ranges of Mask-SSIM [42] on CDF2 [38], i.e., LAA-Net [45], CADDM [21], and SBI [47] for the comparison is directly extracted from [45].
Secondly, in Figure 3b, all models including CNNs and variants of ViTs are trained on FF++ [46] with both real and fake data for $50$ epochs. Following conventional spitting [46], we uniformly extract $128$ and $32$ frames of each video for training and validation, respectively. Hence, there are in total of $460800$ and $22400$ frames for the corresponding task. The weights of models are initialized by pretrained on ImageNet [17]. We employ Adam for CNNs and AdamW for ViTs as optimizers. The learning rate is initially set to $10^{-4}$ and linearly decays to 00 at the end of the training period. All experiments are carried out using a NVIDIA A100 GPU.

### v2

@@ 3 Vision Transformer and Deepfake Detection: Where is the Gap?
Performance with respect to the quality of deepfakes. Here, our goal is to quantify the performance of ViTs and CNNs with respect to the quality of the encountered deepfakes. To that aim, we compare in Figure 3(a) the performance of CNN-based methods (LAA-Net [47], CADDM [21], and SBI [49]) to transformer-based approaches (Plain ViT [23] and Swin [42]) [^2] Note that the small(-S) architecture configuration is employed for both ViT and Swin. The FLOPs and the number of parameters are reported in Section 5. across different Mask Self-Similarity Index Measure (SSIM) [44] ranges on the Celeb-DF [40] dataset. The Mask SSIM has been commonly used for assessing the quality of deepfakes [40, 57, 47] and is obtained by calculating the SSIM [60] score in the facial region between the fake and the corresponding original images. Higher Mask SSIM values indicate a higher perceptual quality of deepfakes [40, 57, 5, 47], generally implying the presence of more localized artifacts. It can be noted from Figure 3(a) that the performance of ViT drops more significantly for higher Mask SSIM values as compared to CNN-based methods. The performance of Swin, on the other hand, does not deteriorate as much due to its ability to extract low-level features with its local window design; however, it still shows less stability than LAA-Net. These observations support our hypothesis. We specify that for a fair comparison with the state-of-the-art (SOTA), we train ViT and Swin using the same data synthesis [49] used in SBI and LAA-Net. More experiments will be provided in supplementary materials.
@@ 4.3.3 Learning-based Local Attention Module (L2-Att)
Our hypothesis is that since the patch size can be too large relative to the area of artifacts, the features encoded within a patch embedding may hold insufficient information about them. Consequently, the implicit self-attention mechanism might overlook or miss important patches, as those containing forgeries can appear too similar to those without. Therefore, we propose an explicit attention mechanism to ensure that the model pays more attention to these critical patches. To this end, by predicting the locations of vulnerable patches, L2-Att diffuses artifact-prone local information through the ViT. In particular, L2-Att first takes the patch embeddings $\mathbf{z}^{L^{\prime}}_{\neg 0}\in\mathbb{R}^{N\times D}$ as input and processes them to produce spatial features as follows,
@@
Method Training set Test set (%)
Real Fake CDF1 CDF2 DFW DFD DFDCP DFDC
@@
Xception [48] $\checkmark$ $\checkmark$ 58.81 65.59 61.18 66.93 65.29 55.37 89.75 85.48 69.90 91.98 45.60 -
FaceXRay [39] $\checkmark$ $\checkmark$ 80.58 73.33 - - - - 95.40 93.34 80.92 72.65 - -
LocalRL [13] $\checkmark$ $\checkmark$ - - 78.26 - - - 89.24 - 76.53 - - -
Multi-attentional [67] $\checkmark$ $\checkmark$ 69.14 74.03 68.26 75.25 73.56 73.79 92.95 96.51 83.81 96.52 - -
RECCE [7] $\checkmark$ $\checkmark$ 49.96 63.04 70.93 70.35 68.16 54.41 98.26 79.42 80.93 92.76 - -
SFDG [59] $\checkmark$ $\checkmark$ - - 75.83 - 69.27 - 88.00 - 73.63 - - -
EIC+IIE [32] $\checkmark$ $\checkmark$ - - 83.80 - - - 93.92 - 81.23 - - -
CADDM [21] $\checkmark$ $\checkmark$ 89.36 93.25 93.88 91.12 74.48 75.23 99.03 99.59 - - 73.85 -
UCF [64] $\checkmark$ $\checkmark$ - - 82.4 - - - 94.5 - 80.5 - - -
M2TR [57] $\checkmark$ $\checkmark$ - - 68.2 - - - - - - - - -
DFDT [35] $\checkmark$ $\checkmark$ - - 88.3 - - - - - 76.1 - - -
TALL-Swin [63] $\checkmark$ $\checkmark$ - - 90.79 - - - - - 76.78 - - -
LSDA [65] $\checkmark$ $\checkmark$ - - 91.10 - - - - - 77.00 - - -
ICT [22] $\checkmark$ 81.43 - 85.71 - - - 84.13 - - - - -
SBI [49] $\checkmark$ 92.53 79.91 93.18 85.16 67.47 55.87 97.56 92.79 86.15 93.24 72.42 -
FakeFormer $\checkmark$ 97.25 98.36 94.45 97.15 81.74 83.72 96.12 98.31 96.30 99.50 78.91 80.01
Table 3: Cross-dataset evaluation. Comparisons in terms of AUC (%) and AP (%) on six benchmark datasets including CDF1 [40], CDF2 [40], DFW [68], DFD [24], DFDCP [19], and DFDC [20]. The results for comparisons are directly extracted from the original papers.
@@ 5.1 Setups
Method Params FLOPs Test set AUC(%)
@@
Method Real Fake Saturation Contrast Block Noise Blur Avg
FaceXray [39] $\checkmark$ $\checkmark$ 97.6 88.5 99.1 49.8 63.8 79.8
LipForensics [28] $\checkmark$ $\checkmark$ 99.9 99.6 87.4 73.8 96.1 91.4
LAA-Net [47] $\checkmark$ 99.9 99.9 99.9 53.9 98.2 90.4
FakeFormer $\checkmark$ 98.1 96.0 97.0 75.3 87.4 90.8
FakeSwin $\checkmark$ 99.8 99.8 99.9 81.6 89.8 94.2
Table 6: Robustness to unseen perturbations. Focusing on vulnerable patches, FakeFormer shows improved robustness to common perturbations, especially to noise.
@@ 5.2 Comparison with State-of-the-art
Comparison with Blending-based Synthesis Approaches. Table 4 and Table 5 compare our method to SOTA when using Blending-Images (BI) [39] and Self-Blended-Images (SBI) [49], respectively. As observed, we generally achieve better performances than SOTA methods using the same blending algorithm on CDF2, DFDC, and DFDCP datasets, despite training on a drastically lower number of data as compared to other ViT-based approaches such as ICT [22]. Moreover, LAA-Net [47], as a CNN-based method, requires a larger model size as compared to our lightweight FakeFormer.
@@
Robustness to Unseen Perturbations. Since deepfakes can be altered by common perturbations on social media. Following the evaluation setting of  [33], experiments evaluating the robustness of FakeFormer against five types of perturbation on FF++ [48] are reported in Table 6. We can observe the superiority of FakeFormer and FakeSwin as compared to existing methods.
@@ 5.3 Ablation Study
Analysis of Parameter $\lambda$. The weight $\lambda$ defined in Eq. (9) is set empirically to $10$ as it yields the best performance on average. We report the results using different values of $\lambda$ within FakeFormer in Table 9. It can be observed the generalization across four testing benchmarks remains robust regardless of the value of $\lambda$.
@@ 8.1 Code Release
Method Mask SSIM Ranges
$0.6$-$0.8$ $0.8$-$0.9$ $0.9$-$0.925$ $0.925$-$0.95$ $0.95$-$0.97$ $0.97$-$1.0$uparrow
ViT+SBI [49] 90.5 94.0 89.2 84.7 82.1 78.0
EFNB4+SBI [49] 89.3 84.1 80.3 83.6 82.8 78.6
Xception+SBI [49] 82.9 80.6 80.5 85.0 82.3 79.4
CADDM [21] 84.1 85.9 82.8 85.4 84.0 81.9
Swin+SBI [49] 90.0 91.0 89.5 85.6 85.8 85.7
LAA-Net [47] 96.1 94.3 92.9 86.0 93.4 91.1
Table 10: Performance analyses across various levels of deepfake quality. AUC (%) comparisons between CNN-based methods (EFNB4 [51]+SBI [49], Xception [14]+SBI [49], CADDM [21], LAA-Net+SBI [47]) and ViT-based methods (ViT [23]+SBI [49], Swin [42]+SBI [49]) across different deepfake quality levels assessed by the Mask SSIM [44] metric. The model is trained on FF++ [48] and evaluated on CDF2 [40].
Input&Patch Size FF++ CDF1 CDF2 DFD DFW DFDCP DFDC
$112$P16161616 81.83 87.93 85.26 73.56 76.87 93.28 72.53
$112$P8888 97.67 97.25 94.45 96.12 81.74 96.30 78.91
$224$P8888 99.93 99.00 96.84 99.54 82.11 96.99 79.01
Table 11: Effect of patch size. AUC (%) comparisons are presented for different setups of input resolution and patch size. The model is trained on FF++ [48] and evaluated on other datasets [40, 24, 68, 20, 19].
@@ 8.2 More Details regarding the Datasets
Datasets. FF++ contains $1000$ original videos and $4000$ deepfake videos. The types of deepfakes include Deepfakes (DF) [50], Face2Face (F2F) [54], FaceSwap (FS) [37], and NeuralTextures (NT) [53]. Moreover, while CDF and DFD represent datasets including high-quality deepfakes [40, 47], DFW stands out as one of the most challenging benchmarks. DFW is collected entirely from the internet without any prior knowledge regarding the used manipulation techniques. Additionally, it includes images with occlusions, illumination changes, different head poses, and diverse backgrounds. Meanwhile, DFDC, including its preview version called DFDCP, comprises more than $128,000$ videos and is therefore one of the largest datasets in the field of deepfake detection.
Data Processing. The RetinaFace [18] algorithm is employed to extract facial bounding boxes. These boxes are enlarged by a factor of $1.25$ and then resized to a specific resolution $(H,W)$ to match the different backbone configurations. To generate pseudo-fakes, the two images are first aligned. For this purpose, the Dlib [36] library is utilized to extract $68$ and $81$ facial landmarks for the two considered blending synthesis algorithms, BI [39] and SBI [49], respectively. Finally, the processed pseudo-fakes are generated for each training iteration.
8.3 Architecture Details
We describe in detail the hyperparameters of the two considered FakeFormer variants as follows:
@@ 8.3 Architecture Details
where the MLP size represents the dimension of hidden layers in MLP, the No. Heads denotes the number of heads in MHSA, Params is the number of parameters, and FLOPs represents the computational cost in terms of floating point operations per second. For FakeSwin architecture, we adopt these two backbone variants from Swin [42], namely:
@@
where $M$ is the window size, $d$ is the query dimension of each head, the expansion layer of each MLP is $\alpha$, and $C_{h}$ denotes the channel number in the hidden layers during the first stage.
8.4 More Details regarding the training Setup in Section 3
In this section, we present more details related to the experimental settings in Section 3 of the main paper.
Specifically, in Figure 3-a, we train vanilla ViT [23] and Swin [42] with SBI data synthesis [49] using the same training settings described in [47, 49]. In addition, the detection performance of three CNN methods in different ranges of Mask-SSIM [44] on CDF2 [40], i.e., LAA-Net [47], CADDM [21], and SBI [49] for the comparison is directly extracted from [47]. Note that, in [49], the authors focus solely on data synthesis. Hence, the used architecture is a standard CNN (EfficientNet-B4 (EFNB4) [51]).
Secondly, in Figure. 3-b, all models including CNNs and variants of ViTs are trained on FF++ [48] with both real and fake data for $50$ epochs. Following conventional spitting [48], we uniformly extract $128$ and $32$ frames of each video for training and validation, respectively. Hence, there are in total of $460800$ and $22400$ frames for the corresponding task. The weights of models are initialized by pretrained on ImageNet [17]. We employ different optimizers as Adam is often used with CNNs [11, 12, 21, 39, 67, 59] and AdamW with ViTs [42, 15, 55, 58]. The learning rate is initially set to $10^{-4}$ and linearly decays to 00 at the end of the training period. All experiments are carried out using a NVIDIA A100 GPU.
8.5 Additional Results and Analysis related to Section 3
In the main paper, Figure. 3-a shows the performance comparison of transformer-based approaches (ViT [23]+SBI [49] and Swin [42]+SBI [49]) and CNN-based methods (EFNB4+SBI [49], CADDM [21], LAA-Net [47]+SBI [49]) across different ranges of Mask-SSIM [44]. Even though the three considered CNN methods are based on EFNB4 [51], most of these methods have been specifically tailored for deepfake detection. Therefore, to support our claims. in addition to EFNB4 [51], we further examine the performance of XceptionNet [14]+SBI [49] on those Mask SSIM ranges under the same comparison setting. Besides EFNB4, Xception is one of the most commonly used backbones in deepfake detection [48, 49, 11, 4, 64]. The overall comparison including the results reported in Figure. 3-b is given in Table. 10. It is interesting to note that Xception performs less effectively than ViT for the first three Mask SSIM range values; however, this changes when considering the last ranges, indicating the presence of high-quality deepfakes with more localized artifacts. The same pattern is also observed for EFNB4 [51]+SBI [49] and other CNN methods, which further confirms our hypothesis (common CNNs such as [51, 14] can capture subtle artifacts more effectively than plain ViT).
8.6 Effect of Patch Size
To further investigate the impact of the patch size on the performance, we provide, in addition to the experiment reported in Figure. 3-b of the main manuscript, the obtained cross-evaluation results of FakeFormer when varying the input size and the patch size in Table. 11. It can be observed that either reducing the patch size or increasing the input resolution contributes to improving the generalization performance of the model. This confirms that the patch size and the input size implicitly affect the extraction of localized features that characterize deepfakes.